OpenStudy (anonymous):

I have two more questions for you @Compassionate First one is: 3x^2+10x=8. AND: 1-squareroot2/2sqareroot3-squareroot6

OpenStudy (compassionate):

Factoring or solving for x?

OpenStudy (anonymous):

first one is factoring. The second one is rationalize.

OpenStudy (compassionate):

I can't really factor the first one. I can find two points on a graph?

OpenStudy (anonymous):

haha.. i do't know what she wants then! I'll just do the second one. Do you know what the answer is for that one?

OpenStudy (compassionate):

For the first one \[x = -4, \frac{2}{3}\] The second one is: \[\frac{1 - \sqrt{2}}{2\sqrt{3} - \sqrt{6}}\] Is that it?

OpenStudy (anonymous):

yes!

OpenStudy (compassionate):

\[- \frac{\sqrt{6}}{6}\]

OpenStudy (anonymous):

thank you sooo much! you're a life saver!

OpenStudy (compassionate):

That's the simplified version... Thanks :)

OpenStudy (compassionate):

Remember, for the first one, x = 4 AND 2/3

OpenStudy (compassionate):

-4"

OpenStudy (anonymous):

yep got it! thanks!

OpenStudy (compassionate):

Going to bed now. Night.